OpenStudy (anonymous):

what is the slope of the line passing through the points (-1,11) and (5,-7)?

OpenStudy (anonymous):

y2-y1/x2-x1= -7-11/5-(-1)=-4/6. The slope is -18/6

OpenStudy (tkhunny):

Let's remember our Order of Operations, please. y2-y1/x2-x1 = \(y2 - \dfrac{y1}{x2} - x1\) and is not as intended. (y2-y1)/(x2-x1) = \(\dfrac{y2-y1}{x2-x1}\) and we have the correct expression. You did it okay in the end, it just isn't right along the way.